# ReDiFine: Reusable Diffusion Finetuning for Mitigating Degradation in the Chain of Diffusion

### v1

@@ A.1 Hyperparameters
We use Stable Diffusion v1.5 [3] as a pretrained text-to-image diffusion model and apply LoRA [21] to fine-tune it at each iteration. We utilize ft-MSE [39] as a fixed VAE to project images into latent space. We augment images with horizontal flip, and no additional augmentation is used. Our implementation is based on [26], with the environment built on PyTorch v2.2.2 [40] and its accompanying torchvision 0.17.2, utilizing CUDA 12.4 on NVIDIA A-100 and L40S GPUs. All default hyperparameters used throughout our paper can be found in Table 2.
@@ A.2 Datasets
We use four image datasets to demonstrate that the degradation occurs universally: Pokemon [22], Kumapi [23], Butterfly [24], and CelebA-1k [25] to investigate various domains including animation, handwriting, and real pictures. All images are resized to $512\times 512$ pixels. BLIP captioner [41] and Waifu Diffusion v1.4 tagger [42] are used to generate text prompts. Samples images and prompts can be found in Table 3.
A.2.1 Pokemon
The Pokemon dataset [22] contains $1008$ images corresponding to the Pokemon indexing from No.1 to No.1008. Prompts are constructed by concatenating captions from Waifu Diffusion tagger and BLIP captioner. We restrict the length of BLIP captions between $50$ and $75$.
Table 3: Samples images and prompts for Pokemon, CelebA-1k, Kumapi, and Butterfly datasets.
@@ A.2.1 Pokemon
A.2.2 CelebA-1k
We subsample the first $1000$ images from CelebA dataset [25] to construct CelebA-1k dataset. Prompts are generated using BLIP captioner with length between $25$ and $50$.
A.2.3 Kumapi
The Kumapi dataset [23] consists of $391$ images with handwriting styles. Prompts are generated using Waifu tagger, and we manually modify some prompts that do not reflect the images well.
A.2.4 Butterfly
Butterfly dataset [24] consists of $832$ images with $8$ species categories. We combine captions generated using the BLIP captioner and descriptions for species from the dataset. The length of captions generated by the BLIP captioner is between $5$ and $75$.
@@ Appendix B Different CFG Scales
This section presents images from Chains of Diffusion for different CFG scales on four datasets outlined above. Figure 10, 11, 12, and 13 are for Pokemon, CelebA-1k, Kumapi, and Butterfly datasets, respectively. For each dataset, results for $5$ different CFG scales are provided to cover high, low, and medium CFG scales. We found $2.5$ to be the optimal medium CFG scale for Pokemon and Kumapi datasets, and $1.5$ for CelebA-1k and Butterfly datasets. With CFG scales lower than the optimal value, low-frequency degradations are observed throughout the four datasets we demonstrate. On the other hand, high-frequency degradations with saturated color and repetitive patterns are universally observed with CFG scales higher than the optimal value. Notably, the optimal value for Pokemon and Kumapi causes the Chain of Diffusion for CelebA-1k and Butterfly datasets to degrade severely.
Figure 10: Chain of Diffusion with different CFG scales for Pokemon dataset. CFG scale $2.5$ is the optimal value we found.
Figure 11: Chain of Diffusion with different CFG scales for CelebA-1k dataset. CFG scale $1.5$ is the optimal value we found.
Figure 12: Chain of Diffusion with different CFG scales for Kumapi dataset. CFG scale $2.5$ is the optimal value we found.
Figure 13: Chain of Diffusion with different CFG scales for Butterfly dataset. CFG scale $1.5$ is the optimal value we found.
Appendix C Hyperparameter investigations to unveil the most impactful factor of degradation
In this section, we present experiment results to reveal the most impactful factor of degradation observed in the Chain of Diffusion. We conduct extensive investigations to check the effects of each hyperparameter outlined in Table 1. Based on the default hyperparameter setting in Table 2, we change each candidate hyperparameter and investigate whether any of them can meaningfully mitigate or affect the degradation. CFG scale is $7.5$ for all cases in this section.
@@ C.1 Training set size
We used CelebA-1k dataset to investigate how the size of the training set (both $D_{0}$ and $D_{k}$) affects the degradation observed in the Chain of Diffusion. We chose this dataset as we can increase or decrease the training set size easily by subsampling from a larger original dataset. We changed the size of the training set to $100$, $250$, $500$, and $2000$. Degradations are observed regardless of the size of the training set, as shown in Figure 14. Degradations occur at earlier iterations when a smaller number of images are used, but it is hardly distinguished in $6$th iteration. We fix the number of parameter updates.
Figure 14: Chain of Diffusion with varying training set size on CelebA-1k dataset. Degradations begin to occur faster with smaller sizes, but all of them results in unrecognizable degree of degradations in $6$th iteration.
@@ C.2 Number of images per prompt
Generating more than one image per prompt to increase the diversity of the training set is a straightforward method to mitigate the degradation in the Chain of Diffusion. We conducted an experiment to investigate how increasing the training set size by generating more images affects the degradation. Figure 15 demonstrates that $5$ times larger images cannot mitigate the degradation. It seems the degradation is slowed down by one iteration ($6$th iteration for the default case and $5$th iteration for the target case look similar), but at the cost of high computations, which makes it an infeasible solution.
Figure 15: Chain of Diffusion when more than one images are generated per prompt on Pokemon dataset. Increasing the training set size by generating more than one image per prompt cannot mitigate the degradation.
@@ C.3 Mixing real images to synthetic sets
Many previous works suggested augmenting the training set with real images to mitigate the degradation observed in iterative training. We investigated if mixing images from the original training set to a synthetic set at each iteration can relieve the degradation. Images are randomly sampled at each iteration and replaced by corresponding real images in the original training set. Figure 16 and 17 demonstrate how the degradation in the Chain of Diffusion differs when $50\%$ and $90\%$ images at each iteration are replaced for Pokemon and CelebA-1k datasets, respectively. It is noteworthy that $10\%$ is far enough to cause the degradation, and $50\%$ rarely slows down the degradation. It is interesting that CelebA-1k dataset is much more vulnerable to degradation.
Figure 16: Chain of Diffusion when real images are randomly sampled and replace synthetic images in every iteration on Pokemon dataset. $50\%$ of real images rarely slow down the degradation, and $10\%$ is enough to commence the degradation.
Figure 17: Chain of Diffusion when real images are randomly sampled and replace synthetic images in every iteration on CelebA-10k dataset. The degradation is rarely slowed down even when only $10\%$ of synthetic images are used for finetuning. Natural images undergo more serious degradation compared to animation images in Figure 16.
@@ C.4 Prompt Set
We hypothesized that the descriptiveness of prompts can affect the degradation of the Chain of Diffusion. Accordingly, we examined different sets of prompts for Pokemon and CelebA-1k datasets. The default prompt set used for Pokemon dataset is the concatenations of Waifu and BLIP captions. BLIP captions are generated to be between $50$ and $75$. We investigated how Waifu prompt set alone and different lengths of BLIP captions can affect the degradation in Figure 18. It is interesting that slightly different styles of high-frequency degradation are observed and short prompt set can relieve repetitive patterns at the cost of decreased diversity. Different lengths of BLIP captions in CelebA-1k dataset demonstrate similar degrees of degradation, but insufficient or excessive descriptiveness of prompt makes images to be different from the original ones as shown in Figure 19.
Figure 18: Chain of Diffusion when different prompts are used on Pokemon dataset. The default prompts are concatenations of BLIP caption of length between $50$ and $75$, and Waifu caption. We compare the Chain of Diffusion with Waifu captions, BLIP captions (used for default prompts), and short BLIP captions. Short captions are generated to have lengths smaller than $25$.
Figure 19: Chain of Diffusion when different prompts are used on CelebA-1k dataset. The default prompts have lengths between $25$ and $50$. We compare the Chain of Diffusion with the lengths of prompts larger than $50$ (Long) and smaller than $25$ (Short).
@@ C.5 U-Net and Text-Encoder
Figure 20 shows the Chains of Diffusion when one of U-Net or text encoder is finetuned. Similar degradation happens when the text encoder is not updated (second row), and the degradation pattern is different when the U-Net of Stable Diffusion is not updated since image qualities that models can generate do not change in this case. However the contents of images are not preserved as text encoder is updated.
Figure 20: Chain of Diffusion when one of U-net and text encoder is finetuned on Pokemon dataset.
@@ C.6 Number of diffusion steps
We examined if the number of diffusion steps during generation is insufficient and causes the degradation. Figure 21 demonstrates that the number of diffusion steps cannot improve the Chain.
@@ C.7 Number of Epochs
We also tested if the training is insufficient or excessive for our default setting. However, images from the first iterations look similar in quality, as shown in Figure 21, and the Chains result in similar degradations. We finetune models for $100$ epochs as the default setting since loss values still decrease in $50$ epochs.
Figure 21: Chain of Diffusion when the number of diffusion steps and training epochs differ on Pokemon dataset. Both more diffusion steps and different training epochs cannot mitigate the degradation, and similar patterns are observed.
@@ C.8 Learning Rate
Similarly, we examined in Figure 22 if finetuning is properly done with our default setting. We multiplied the learning rate for U-Net and text encoder by $10$ and $0.1$. Images of the first iterations indicate that the default values are appropriate for finetuning, and the degradation constantly occurs for different learning rate.
Figure 22: Chain of Diffusion when learning rate differs and Gaussian noise is added to the original training set on Pokemon dataset.
@@ C.9 CLIP Skip
We examined if CLIP skip hyperparameter can affect the degradation. CLIP skip determines what intermediate feature from CLIP text encoder is used as text embedding for conditional generation. Smaller value means the feature closer to the output of CLIP text encoder is used, and larger values for features closer to the input text. Figure 23 shows that this hyperparameter rarely changes the degradation patterns.
Figure 23: Chain of Diffusion when different CLIP skip hyperparameter is used on Pokemon dataset. CLIP skip hyperparameter has negligible effect to the degradation.
@@ C.10 Adding Gaussian noise to the original training set
We inspected if there are any differences between real and synthetic images, and if they can be cause the degradation. Random Gaussian noise is added to the original training set $D_{0}$. Figure 22 shows that the characteristics of the original training set have some effects, but the degradation still happens. This aligns with our observations that the degradations are universal to real, animation, and hand-written images.
@@ C.11 Stable Diffusion XL
We investigated how image degradation occurs for different Stable Diffusion models. We note here that optimal hyperparameters for finetuning SDXL using LoRA have not been investigated enough yet, so we use the same hyperparameters to Stable Diffusion v1.5 and it can be suboptimal. Due to the space complexity, we reduce batch size to $2$. We also use the same resolution $512\times 512$ for this model as the first iteration images are impressively high-quality.
Figure 24: Chain of Diffusion of SDXL on Pokemon dataset. Optimal hyperparameters for finetuning SDXL have not been investigated enough yet, so our experiments are largely based on hyperparameters for Stable Diffusion v1.5. We note that images for Pikachu are provided to align with other cases, but SDXL has a huge bias in generating Pikachu well while it fails for other indices. Nonetheless, there are clear changes with different CFG scales.
Appendix D ReDiFine

### v2

@@ A.1 Hyperparameters
We finetune Stable Diffusion v1.5 [24] using LoRA [25] at each iteration, with ft-MSE [48] as a fixed VAE to project images into latent space. Horizontal flip is the only image augmentation applied. Our implementation follows [26] and is built on PyTorch v2.2.2 [49], with torchvision 0.17.2, running on CUDA 12.4 using NVIDIA A-100 and L40S GPUs. All default hyperparameters are listed in Table 1.
@@ A.2 Datasets
We use four image datasets to demonstrate the universal nature of degradation: Pokemon [27], Kumapi [28], Butterfly [29], and CelebA-1k [30], covering animation, handwriting, and real images. All images are resized to $512\times 512$ pixels. Text prompts are generated using BLIP captioner [50] and Waifu Diffusion v1.4 tagger [51]. Sample images and prompts can listed in Table 2.
Pokemon.
The Pokemon dataset [27] contains $1008$ images indexed by number, with prompts combining BLIP captions (length $50$-$75$ words) and Waifu Diffusion tagger.
Table 2: Sample images and prompts for Pokemon, CelebA-1k, Kumapi, and Butterfly datasets.
@@
CelebA-1k.
CelebA-1k is a subsample of $1000$ images from CelebA [30], with BLIP captions ($25$-$50$ words).
Kumapi.
Kumapi [28] includes $391$ handwriting-style images, with Waifu-generated prompts and manual adjustments.
Butterfly.
Butterfly [29] includes $832$ images across $8$ species, using BLIP captions (length $50$-$75$ words) and species descriptions.
@@ Appendix B Different CFG Scales
This section presents images from Chain of Diffusion at various CFG scales for the four datasets. Figure 8, 9, 10, and 11 corresponds to Pokemon, CelebA-1k, Kumapi, and Butterfly, respectively. Each dataset shows results for five CFG scales, covering high, medium, and low values. The optimal medium CFG scale is $2.5$ for Pokemon and Kumapi, and $1.5$ for CelebA-1k and Butterfly. Lower-than-optimal CFG scales lead to low-frequency degradations, while higher-than-optimal scales result in high-frequency degradations with saturated colors and repetitive patterns. Notably, the optimal CFG for Pokemon and Kumapi causes severe degradation in CelebA-1k and Butterfly.
Figure 8: Chain of Diffusion for Pokemon at various CFG scales. The optimal CFG scale is $2.5$.
Figure 9: Chain of Diffusion for CelebA-1k at various CFG scales. The optimal CFG scale is $1.5$.
Figure 10: Chain of Diffusion for Kumapi at various CFG scales. The optimal CFG scale is $2.5$.
Figure 11: Chain of Diffusion for Butterfly at various CFG scales. The optimal CFG scale is $1.5$.
Appendix C Hyperparameter investigations to unveil the most significant factor of degradation
This section presents experimental results identifying the most significant factors contributing to degradation in the Chain of Diffusion. We systematically vary each hyperparameter from Figure 2, using the default settings from Table 1, to assess their impact on degradation. The CFG scale is fixed at $7.5$ for all cases.
@@ C.1 Training set size
We used CelebA dataset [30] to examine how training set size (both $D_{0}$ and $D_{k}$) impacts degradation in the Chain of Diffusion. By subsampling, we adjusted the training set to $100$, $250$, $500$, and $2000$ images. Degradation occurs regardless of dataset size, as shown in Figure 12, but appears earlier with smaller sets. By the $6$th iteration, images degrade severely for all cases. The number of parameter updates was kept constant across all dataset sizes.
Figure 12: Chain of Diffusion on CelebA dataset with varying training set sizes. Degradation occurs faster with smaller sets, but all result in severe degradation by the $6$th iteration.
@@ C.2 Number of images per prompt
Generating multiple images per prompt is a simple way to increase training set diversity and can be considered as a solution to mitigate degradation in the Chain of Diffusion. We tested this by generating $5$ times more images. As shown in Figure 13, while degradation is slightly delayed (by one iteration), it remains unmitigated, and the high computational cost makes this approach impractical.
Figure 13: Chain of Diffusion on Pokemon dataset with multiple images generated per prompt. Increasing the training set size in this way does not mitigate degradation.
@@ C.3 Mixing real images to synthetic sets
Many previous works suggest augmenting the training set with real images to mitigate degradation during iterative training. We investigated whether mixing images from the original training set into the synthetic set at each iteration could alleviate this issue. At each iteration, images are randomly replaced with corresponding real images. Figure 14 and 15 show how degradation in the Chain of Diffusion varies when $50$% and $90$% of images are replaced for Pokemon and CelebA-1k datasets, respectively. Notably, even $10$% synthetic images are sufficient to induce degradation, and $50$% replacement rarely slows it down. CelebA-1k dataset appears to be significantly more susceptible to degradation.
Figure 14: Chain of Diffusion on Pokemon dataset with real images randomly replacing synthetic images at each iteration. A $50$% replacement rarely slows degradation, while $10$% synthetic images are sufficient to initiate it.
Figure 15: Chain of Diffusion on CelebA-1k dataset with real images randomly replacing synthetic images at each iteration. A $50$% replacement rarely slows degradation, while $10$% synthetic images are sufficient to initiate it. It suffers from more severe degradation than Pokemon dataset as compared with Figure 14.
@@ C.4 Prompt Set
We hypothesized that the descriptiveness of prompts influences degradation in the Chain of Diffusion. We tested various prompt sets for Pokemon and CelebA-1k datasets. For Pokemon dataset, the default prompt set consists of concatenated Waifu and BLIP captions, with BLIP captions ranging from $50$ to $75$ words. Figure 16 shows how using only Waifu prompts and varying BLIP caption lengths affect degradation. Notably, different styles of high-frequency degradation were observed; shorter prompts reduced repetitive patterns but decreased diversity. In CelebA-1k dataset, varying BLIP caption lengths resulted in similar degradation levels, but prompts that were either insufficiently or excessively descriptive caused the images to deviate from the originals, as shown in Figure 17.
Figure 16: Chain of Diffusion on Pokemon dataset with different prompts. The default prompts are concatenations of BLIP captions ($50$-$75$ words) and Waifu captions. We compare the Chain of Diffusion using default captions, Waifu captions, short (less than $25$ words) and long ($50$-$75$ words) BLIP captions.
Figure 17: Chain of Diffusion on CelebA-1k dataset with different prompts. The default prompts range from $25$ to $50$ words. We compare the Chain of Diffusion using longer prompts (over $50$ words) and shorter prompts (under $25$ words).
@@ C.5 U-Net and Text-Encoder
Figure 18 illustrates the Chain of Diffusion with either the U-Net or text encoder finetuned. When the text encoder is not updated (second row), similar degradation occurs. However, the degradation pattern changes when the U-Net is not updated, as the model’s ability to generate images remains unchanged. In contrast, updating the text encoder results in a loss of image content preservation.
Figure 18: Chain of Diffusion on Pokemon dataset with either the U-Net or text encoder finetuned.
@@ C.6 Number of diffusion steps
We investigated whether an insufficient number of diffusion steps during generation contributes to degradation. Figure 19 shows that increasing the number of diffusion steps does not enhance the Chain of Diffusion.
@@ C.7 Number of Epochs
We also examined whether insufficient or excessive training affects our default setting. As shown in Figure 19, images from the initial iterations exhibit similar quality, resulting in comparable degradations. We set the default finetuning to $100$ epochs since loss values continue to decrease after $50$ epochs.
Figure 19: Chain of Diffusion on Pokemon dataset with varying diffusion steps and training epochs. Both increased diffusion steps and differing training epochs fail to mitigate degradation, resulting in similar patterns.
@@ C.8 Learning Rate
Similarly, we assessed finetuning adequacy in Figure 20 by adjusting the learning rates for the U-Net and text encoder by x10101010 and x0.10.10.10.1. Images from the initial iterations show that the default values are suitable for finetuning. Although the styles are different, degradation consistently occurs across different learning rates.
Figure 20: Chain of Diffusion on Pokemon dataset with varying learning rates and added Gaussian noise to the original training set.
@@ C.9 CLIP Skip
We investigated the impact of the CLIP skip hyperparameter on degradation. The CLIP skip determines which intermediate feature from the CLIP text encoder is used as the text embedding for conditional generation, with smaller values selecting features closer to the output and larger values selecting those nearer to the input text. As shown in Figure 21, this hyperparameter has minimal effect on degradation patterns.
Figure 21: Chain of Diffusion on Pokemon dataset using different CLIP skip hyperparameters. The CLIP skip hyperparameter shows a negligible effect on degradation.
@@ C.10 Adding Gaussian noise to the original training set
We examined whether differences between real and synthetic images contribute to degradation by adding random Gaussian noise to the original training set $D_{0}$. Figure 20 illustrates that while the characteristics of the original training set have some effect, degradation still occurs. This supports our findings that degradations are universal across real, animation, and handwritten images.
@@ C.11 Stable Diffusion XL
We investigated image degradation for a different Stable Diffusion model, noting that the optimal hyperparameters for finetuning SDXL using LoRA are not well established. Consequently, we applied the same hyperparameters used for Stable Diffusion v1.5, which may be suboptimal. To manage space complexity, we reduced the batch size to $2$ and maintained a resolution of $512\times 512$, as the first iteration images exhibit impressive quality. Results are presented in Figure 22.
Figure 22: Chain of Diffusion of SDXL on Pokemon dataset. Due to insufficient investigation into optimal hyperparameters for finetuning SDXL, our experiments largely rely on those from Stable Diffusion v1.5.
Appendix D Quantitative Trait Modeling
D.1 Proof of Theorem 1
Proof. The difference between the successive means in quantitative trait modeling is given by:
$\Delta\mu_{t}=\mu_{t+1}-\mu_{t}=h_{t}^{2}(\mu_{t}^{\prime}-\mu_{t}).$ (6)
When Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$ is truncated on both sides with $r_{1}$ and $r_{2}$ ratios, the mean and variance of truncated Gaussian distribution are expressed as:
$\mu_{trun}=\mu-\frac{\varphi(\beta)-\varphi(\alpha)}{\Phi(\beta)-\Phi(\alpha)}\sigma,$ (7)
$\sigma_{trun}^{2}=(1-\frac{\beta\varphi(\beta)-\alpha\varphi(\alpha)}{\Phi(\beta)-\Phi(\alpha)}-(\frac{\varphi(\beta)-\varphi(\alpha)}{\Phi(\beta)-\Phi(\alpha)})^{2})\sigma^{2},$ (8)
where $\varphi$ and $\Phi$ are the probability density function (PDF) and cumulative distribution function (CDF) of the standard normal distribution, and $\alpha=\Phi^{-1}(r_{1})$ and $\beta=\Phi^{-1}(1-r_{2})$. Accordingly, given $\mu=\mu_{t}$ and $\sigma^{2}=\sigma_{P,t}^{2}$ with $\mu_{t}^{\prime}=\mu_{trun}$ and $\sigma_{G,t+1}^{2}=\sigma_{P,t}^{\prime 2}=\sigma_{trun}^{2}$, we have
$\Delta\mu_{t}=h_{t}^{2}(\mu_{t}^{\prime}-\mu_{t})=\frac{\sigma_{G,t}^{2}}{\sigma_{P,t}^{2}}c_{1}\sigma_{P,t}=c_{1}\frac{\sigma_{P,t-1}^{\prime 2}}{\sigma_{P,t}^{2}}\sigma_{P,t}=c_{1}c_{2}\frac{\sigma_{P,t-1}^{2}}{\sigma_{P,t}^{2}}\sigma_{P,t},$ (9)
where $c_{1}=|\frac{\varphi(\beta)-\varphi(\alpha)}{\Phi(\beta)-\Phi(\alpha)}|$ and $c_{2}=1-\frac{\beta\varphi(\beta)-\alpha\varphi(\alpha)}{\Phi(\beta)-\Phi(\alpha)}-(\frac{\varphi(\beta)-\varphi(\alpha)}{\Phi(\beta)-\Phi(\alpha)})^{2}$ when $r_{1}>r_{2}$. On the other hand, the mean phenotypes decreases when $r_{1}<r_{2}$ as:
$\Delta\mu_{t}=-c_{1}c_{2}\frac{\sigma_{P,t-1}^{2}}{\sigma_{P,t}^{2}}\sigma_{P,t}.$ (10)
Furthermore, the phenotype variance converges over time as:
$\sigma_{P,t}^{2}=\sigma_{G,t}^{2}+\sigma_{E}^{2}=\sigma_{P,t-1}^{\prime 2}+\sigma_{E}^{2}=c_{2}\sigma_{P,t-1}^{2}+\sigma_{E}^{2}.$ (11)
$\sigma_{P,t}^{2}-\frac{\sigma_{E}^{2}}{1-c_{2}}=c_{2}(\sigma_{P,t-1}^{2}-\frac{\sigma_{E}^{2}}{1-c_{2}})=\cdots=c_{2}^{t}(\sigma_{P,0}^{2}-\frac{\sigma_{E}^{2}}{1-c_{2}}).$ (12)
As a result, the phenotype variance converges to $\frac{\sigma_{E}^{2}}{1-c_{2}}$ for $10<c_{2}<$ 1 (the variance of truncated distribution is smaller than the variance of the original distribution), and the mean asymptotically increases (decreases) by $\frac{c_{1}c_{2}}{\sqrt{1-c_{2}}}\sigma_{E}$ per iteration when $r_{1}>r_{2}$ ($r_{2}>r_{1}$). $\square$
D.2 Simulation setup
Our simulation aims to demonstrate that our experimental results can be modeled using quantitative trait modeling. We show that the radial sum of power spectra of images is one of the phenotypes explained by our theoretical analysis.
D.2.1 Computing the radial sum of power spectra.
The power spectra of images are computed as the square of the magnitude of the 2D Fourier transform. Here, we demonstrate how to compute the radial sum of power spectra in order to use the sum of radial power spectra above a certain threshold. Given a set of images $\{I_{i}\}$ where $i\in\{1,2,\cdots,N\}$, the 2D Discrete Fourier Transform (DFT) of an image $I_{i}$ of size $M\times N$ is computed as:
$F_{i}(u,v)=\sum^{M-1}_{x=0}\sum^{N-1}_{y=0}I_{i}(x,y)e^{-2\pi j(\frac{ux}{M}+\frac{vy}{N})}$ (13)
where $F_{i}(u,v)$ represents the frequency component at coordinates $(u,v)$. The power spectrum of an image $I_{i}$ at $(u,v)$ is the square of the magnitude of its Fourier transform $P_{i}(u,v)=|F_{i}(u,v)|^{2}$. We compute the radial sum of the power spectra using a norm of each frequency component $(u,v)$ as $r(u,v)=\sqrt{(\frac{u}{M})^{2}+(\frac{v}{N})^{2}}$. Given the threshold frequency $\tau$, we compute the radial sum of the high-frequency power spectra of an image $I_{i}$ as
$S_{i}=\sum^{M-1}_{u=0}\sum^{N-1}_{v=0}P_{i}(u,v)\cdot{\rm I}(r(u,v)>\tau)$ (14)
where ${\rm I}(\cdot)$ is an indicator function. We compute the sum of high-frequency components because low-frequency components tend to be noisy and use the threshold frequency of $0.02$. Then, the total sum of power spectra for all images can be written as:
$S=\sum^{N}_{i=0}S_{i}=\sum^{N}_{i=0}\sum^{M-1}_{u=0}\sum^{N-1}_{v=0}P_{i}(u,v)\cdot{\rm I}(r(u,v)>\tau).$ (15)
Practically, we shift the Fourier transform maps so that the frequency norm larger than $\frac{1}{\sqrt{2}}$ is considered to be in the opposite direction. Code for detailed implementation comes from the official code of [45].
Figure 23: Results comparing the simulation for mutations and power spectra of images generated by ReDiFine (with CFG scale $7.5$). Power spectra and simulation are plotted in solid and dotted lines, respectively. Our modifications to heritability and selection process successfully demonstrate changes occur to images by ReDiFine.
D.2.2 Simulation parameters matching different CFG scales.
We simulate different selection strategies using truncation ratios $r_{1}$ and $r_{2}$. To model high, medium, and low CFG scales from our experiments, we apply $(0.025,0.675)$, $(0.5,0.09)$, and $(0.95,0.0002)$ for $(r_{1},r_{2})$, respectively, which correspond to CFG scales of $7.5$, $2.5$, and $1.0$. For initial values in the simulation, we compute the mean ($0.027$) and standard deviation ($0.056$) from the original training set (iteration 00), using them as the initial values for mean and genetic standard deviation for our simulation. We set the environmental standard deviation to $0.25$.
D.2.3 ReDiFine and mutations.
ReDiFine combines condition drop finetuning and CFG scheduling, inspired by the mutation mechanism that compensates for distribution shift and maintains genetic distributions in population genetics. We apply two modifications to our theoretical analysis of Section 4—adding mutation variance to heritability and smoothing truncations—to simulate the effects of ReDiFine in the Chain of Diffusion. Specifically, we add the mutation standard deviation $\sigma_{M}$ of $0.1$ to heritability as:
$h_{t}^{2}=\frac{\sigma_{G,t}^{2}}{\sigma_{P,t}^{2}+\sigma_{M}^{2}}=\frac{\sigma_{G,t}^{2}}{\sigma_{G,t}^{2}+\sigma_{E}^{2}+\sigma_{M}^{2}},$ (16)
representing the randomness added to each iteration due to mutations. This influences the effects of previous iteration to current iteration, which reflect finetuning. Moreover, we apply exponential tails to truncations instead of cut-off thresholds where samples outside the truncation area can be randomly selected with exponential distribution ($e^{-\alpha d(x)}$ where $d(x)$ is a distance to truncation zone and we use $\alpha=0.1$). This modified selection simulates the effect of CFG scheduling during image generation.
Figure 23 shows the effects of two modifications to our simulation results, and they closely align with the power spectra of images generated by ReDiFine. This demonstrates that the effects of ReDiFine can be understood as interference similar to mutations in population genetics. This suggests further research on model collapse motivated from other fields like biology.
Appendix E ReDiFine